A random sample of 20 drinks from a soft-drink machine has an average content of 21.9 deciliters, with a standard deviation of 1.42 deciliters. At 0.05 level of significance, test the hypothesis that μ = 22. 2 against the alternative hypothesis μ < 22.2. Assume that the distribution is normal.


1
Expert's answer
2022-06-07T10:56:41-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=22.2"

"H_1:\mu<22.2"

This corresponds to a left-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=19" and the critical value for a left-tailed test is "t_c =-1.729133."

The rejection region for this left-tailed test is "R = \{t:t<-1.729133\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{21.9-22.2}{1.42/\sqrt{20}}=-0.9448"


Since it is observed that "t=-0.9448>-1.729133=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for left-tailed, "df=19" degrees of freedom, "t=-0.9448" is "p=0.178311," and since "p= 0.178311>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is less than 22.2, at the "\alpha = 0.05" significance level.